Circular No. 7067

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

     URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1998 X1 (ODAS)

The Observatoire de la Côte d’Azur-Deutsches Zentrum für Luft und Raumfahrt Asteroid Survey (ODAS) reports the discovery of a comet.

1998 UT α_{2000} δ_{2000} m_{1} Observer
Dec. 15.17384 9^{h}59^{m}03.^{s}68 +13^{o}54′35′′.1 18.3 Maury
15.21959 9 59 04.53 +13 54 30.4 18.3  _{”}
16.04458 9 59 20.39 +13 53 16.7  Tichý
16.07227 9 59 20.93 +13 53 13.1  _{”}
16.08374 9 59 21.05 +13 53 12.6  Šarounová
16.13874 9 59 22.04 +13 53 08.5 18.9 Maury
16.16045 9 59 22.37 +13 53 05.9 19.1  _{”}
16.17472 9 59 22.70 +13 53 06.4 19.4 Colas
16.26022 9 59 24.10 +13 52 59.1 19.1  _{”}
16.98021 9 59 36.79 +13 51 58.9 18.9 Stoss
17.18946 9 59 39.85 +13 51 43.8  Tichý
17.20978 9 59 40.26 +13 51 42.5  _{”}
17.21033 9 59 40.24 +13 51 43.2 18.6 Colas
17.26252 9 59 41.03 +13 51 39.2 19.4  _{”}

A. Maury, D. Albanese, G. Hahn and M. Hoffmann (Caussols). 0.90-m

Schmidt. Dec. 15: 1′ tail in p.a. 300^{o}. Dec. 16: image diffuse. M. Tichý and Z. Moravec (Kleť). 0.57-m f/5.2 reflector. Dec. 17: image

diffuse; 10′′ tail in p.a. 291^{o}. L. Šarounová (Ondřejov). 0.65-m f/3.6 reflector. Small, faint coma; prob-

able faint and wide tail. F. Colas and J. Lecacheux (Pic du Midi). 1.05-m f/6 reflector. R. Stoss, S. Kluegl and P. Geffert (Heppenheim). 0.45-m f/4.4 reflector.

The orbit of this comet is extremely indeterminate, but the following parabolic elements should provide a useful early ephemeris. The comet is probably a short-period one.

T = 1998 Aug. 19.081 TT  ω = 63.^{o}239
Ω = 7.550 }2000.0
q = 1.87795 AU  i = 1.261
1998/99 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 18 9^{h}59.^{m}87 +13^{o}50.^{′}8 1.760 2.393 118.^{o}6 21.^{o}2 18.0
23 10 00.71 +13 47.1 1.744 2.430 123.5 19.7 18.1
28 10 00.79 +13 47.3 1.731 2.467 128.6 18.2 18.1
Jan. 2 10 00.16 +13 51.0 1.722 2.505 133.8 16.4 18.2
1998 December 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden